OpenStudy (anonymous):

Fan and medal--A square sheet of paper is folded diagonally. Find the length of the diagonal.

OpenStudy (anonymous):

A.5 cm B.10 cm C.5 sqrt. 2 D. 10 sqrt. 2

OpenStudy (kidrah69):

\[a^2+b^2=c^2\]

OpenStudy (kidrah69):

plug in the numbers

OpenStudy (anonymous):

5 5

OpenStudy (anonymous):

so 25+25 = 50 = 25

OpenStudy (kidrah69):

yep but square it

OpenStudy (kidrah69):

where did u get 50=25 from?

OpenStudy (anonymous):

50*50=2500= 50

OpenStudy (anonymous):

it was a mistake

OpenStudy (kidrah69):

oh ok so 50^2 good job :)

OpenStudy (anonymous):

i dont see which of my answer choices i can get tho

OpenStudy (anonymous):

i think D but i really dont know

OpenStudy (kidrah69):

well you didnt put any answer choices so i wouldnt know

OpenStudy (anonymous):

I did, they are with the pic but here :) A.5 cm B.10 cm C.5 sqrt. 2 D. 10 sqrt. 2

OpenStudy (kidrah69):

oops sorry =/ im not sure i think we did something wrong @Data_LG2

OpenStudy (anonymous):

its okay and :( oh man

OpenStudy (kidrah69):

hmmm well we have the Pythagorean theorm \[a^2+b^2=c^2\] we plugged in the numbers and got \[5^2+5^2=c^2\]

OpenStudy (anonymous):

i have other ones you may be able to help with instead

OpenStudy (anonymous):

and true

OpenStudy (kidrah69):

5x5=25 5x5=25 25+25=50

OpenStudy (anonymous):

yes

OpenStudy (kidrah69):

50^2=2500

OpenStudy (anonymous):

50*50=2500

OpenStudy (anonymous):

agreee

OpenStudy (kidrah69):

@amistre64

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

We find the square root of 50..we don't square it.. \(\sqrt{50}\) Then we simplify it

OpenStudy (anonymous):

finally, it opens ! x_x \(\sf c= \sqrt{5^2+5^2}\\c=\sqrt{25+25}\\c=\sqrt{50}\) like iGreen said. simplify it :P